OpenStudy (anonymous):

Two protons are released from rest when they are 0.750 nm apart. What is the maximum speed they will reach? I already know the answer but I don't understand how to derive the correct formula to use to get that answer please help guys

OpenStudy (irishboy123):

\[k\frac{q.q}{r_o} = 2(\frac{1}{2}m \dot r_{\infty}^2)\]??

OpenStudy (anonymous):

how do you get that? @IrishBoy123

OpenStudy (irishboy123):

potential energy of the system is \(\large U = k\frac{q.q}{r} \) when the protons are distance r apart, yes?